10.00grams of magnesium acetate is dissolved in 100.00 ml of DI water . The molar mass is 142.41g/mol . Calculate the molarity
1
Expert's answer
2020-06-16T15:00:22-0400

Solution:

Molarity = mol solute / L of solution


Magnesium acetate = Mg(C2H3O2)2

Moles of Mg(C2H3O2)2 = Mass of Mg(C2H3O2)2 / Molar mass of Mg(C2H3O2)2

Moles of Mg(C2H3O2)2 = (10.00 g) / (142.41 g/mol) = 0.07022 mol

mol solute = 0.07022 mol


L of solution = 100.00 mL = 0.1000 L


Therefore,

Molarity of Mg(C2H3O2)2 = (0.07022 mol) / (0.1000 L) = 0.7022 mol/L = 0.7022 M


Answer: The molarity of magnesium acetate is 0.7022 M.